OpenStudy (anonymous):

What is the slope of the line that passes through the pair of points (2, 9), and (5, 12)?

OpenStudy (anonymous):

Use this equation/formula. This will be your mathematical bible when it comes to solving questions like this. \[y-y1=m(x-x1)\] Where m is the slope and (x1, y1) is any one of the two points you were given. The only part missing is the slope. To find the slope with two given points that both lie on the line, you would then use this equation: \[\frac{ y2-y1 }{ x2-x1 }\]. It can be vice versa whe

OpenStudy (anonymous):

\[m= (12-9)/(5-2)\] \[m= 3/3\] \[m=1\]Now you have the slope. All you have to do is substitute one of the points given with the slope into the first equation/formula I've given you.

OpenStudy (anonymous):

To find slope:\[\frac{ y _{2}- y _{1} }{ x _{2}-x _{1} }\] You have the points: (2, 9), and (5, 12) Where y2= 12, y1= 9 and x2= 5 and x1= 2 Just set up and solve:\[\frac{ 12-9 }{ 5-2}\]